OpenStudy (anonymous):

Solve for x. 3 ^(2x+2) = 8

OpenStudy (anonymous):

Im confused on how to solve this with 2x+2 being a exponent

zepdrix (zepdrix):

\[\huge 3^{\color{royalblue}{2x+2}}=8\] Hmm our X is in the exponent! That's no good. So we'll have to apply some fancy work with logarithms.

zepdrix (zepdrix):

Take the natural log of both sides.\[\huge \ln\left(3^{\color{royalblue}{2x+2}}\right)=\ln 8\]

zepdrix (zepdrix):

An important rule of logarithms states,\[\huge \log (a^b)= b\cdot \log (a)\]We're going to apply this to our problem, giving us,\[\huge \color{royalblue}{(2x+2)}\ln(3)=\ln 8\]Understand what we did there? :o

OpenStudy (anonymous):

so I do the In for 3 and 8?

zepdrix (zepdrix):

We did the natural log of the entire left side, and the entire right side. Then, using a rule of logarithms, we were able to bring the exponent down in front of the log. Now we have just a little bit more work to solve for x.

zepdrix (zepdrix):

Divide both sides by ln(3),\[\huge \color{royalblue}{2x+2}=\frac{\ln8}{\ln3}\] Then we'll subtract 2 from each side,\[\huge \color{royalblue}{2x}=\frac{\ln8}{\ln3}-2\] Then we'll divide both sides by 2,\[\huge \color{royalblue}{x}=\frac{\ln8}{2\ln3}-1\]

zepdrix (zepdrix):

It's a bit of work to solve this one, getting confused on the logarithms?

OpenStudy (anonymous):

I got rounded 0.14

OpenStudy (anonymous):

is that correct and the more complicated ones Im confused with

zepdrix (zepdrix):

Hmm I got -0.0536 Lemme try it again to make sure I didn't make a mistake.

zepdrix (zepdrix):

Yah I'm still coming up with -0.0536. Yah I checked it on wolfram, same answer. Having a little trouble punching it in correctly on the calculator? :o

OpenStudy (anonymous):

I guess so. im doing it on my omputer calculator. i did 8 in / 2*3 in = then subtract by one

zepdrix (zepdrix):

Well I'm not sure what your computer calculator looks like. But on a normal TI calculator it looks something like this. \[\huge ((\ln(8))\div (2 \ln(3)))-1\] I definitely prefer to do it in pieces instead of using this many parentheses.

zepdrix (zepdrix):

Are you using the Windows Calculator?

zepdrix (zepdrix):

For the Windows Calculator you would put it in like this,\[\huge 8 \ln\quad /\quad 2 \quad / \quad 3 \ln \qquad \text{Enter} \qquad - \quad 1\]